OpenStudy (anonymous):

d = 55t, solve for t

OpenStudy (aravindg):

divide by 55 on both sides

OpenStudy (ryan123345):

Divide by 55 to isolate the variable.

OpenStudy (mathstudent55):

t is being multiplied by 55. The opposite operation of multiplication is division. To get rid of 55, divide 55t by 55. When you deal with an equation, you must do the same thing to both sides, so divide both sides by 55. d = 55t d/55 = 55t/55 d/55 = t t = d/55